OpenStudy (anonymous):

a car with a mass of 2,200 kg can accelerate from 0 to 30 m/s in 6 seconds . a. assuming all of the force generated by the engine goes into acceleration of the car ,how much force is the engine producing?

OpenStudy (anonymous):

v=u+at will get you the accelreation. Then use that a in F=ma and you will get the force

OpenStudy (anonymous):

is it 2,200 +30 =

OpenStudy (anonymous):

30=0+6a So a is 5 Then F = ma= 2200*5=11000

OpenStudy (anonymous):

i got 11000 as an answer is that right for this problem ?

OpenStudy (anonymous):

Well done. Yes.